a restaurant offers the possibility of 168 three course dinners. Each dinner has an appetizer, an entree' and a dessert. If the number of appetizers decreases from 7 to 5, find how many fewer possible three-course dinners the restaurant offers.
Answer by richwmiller(17219)

a*b*c=168
7*b*c=168
b*c=168/7=24
now we have
5*24=120
168-120=48 or 2*24
There are 48 fewer possible three course dinners